As team power plant receives heat from a furnace at a rate of 280 GJ/h. Heat losses to the surrounding air from the steam as it passes through the pipes and other components are estimated to be about 8 GJ/h. If the waste heat is transferred to the cooling water at a rate of 145 GJ/h, determine (a) whethupoometoutput and (b) the thermal efficiency of this power plant.

1
Expert's answer
2022-02-25T08:43:02-0500

determine (a) net power output:

"\dot{W}_{out}=\dot{Q}_{in}-\dot{Q}_{out}-\dot{Q}_{loss}"

"\dot{W}_{out}=(280-145-8).(\frac{1000}{3600})=35.278MW"


b) The thermal efficiency of the power plant is:

"\frac{\dot{Q}_{in}-\dot{Q}_{out}-\dot{Q}_{loss}}{\dot{Q}_{in}}\times 100\%" "=\frac{280-145-8}{280}\times 100\%=45.357\%"